OpenStudy (anonymous):

Estimate the volume of material in a cylindrical shell with height 30 in., radius 10 in., and shell thickness 0.5 in.

OpenStudy (anonymous):

therefore radius is 9.5 and height is 29 because of the thickness of the material v = 2(pi)r^2h v = 2 (pi) (9.5)^2 (29)